OpenStudy (anonymous):

Please helpp?! I did the problem I just need to know if im correct. Which of the following represents the area of a rectangle whose length is x – 7 and whose width is x + 12? x2+ 5x – 84 x2– 84 x2– 5x – 84 x2– 84x + 5 I got C

OpenStudy (anonymous):

ur sexy

OpenStudy (anonymous):

Its A not C and really mr. long name really?

OpenStudy (anonymous):

why is it A? :)

OpenStudy (anonymous):

The 5x is a positive because the 12 is positive and you subtract 7 from that

OpenStudy (anonymous):

Yes i also got the C : it is completely correct it will be : x^2 +5x-84

OpenStudy (anonymous):

a bigger posative and a smaller negative lesave a positive

OpenStudy (anonymous):

Its B for sure...

OpenStudy (anonymous):

ya but c is X^2 !!-!! 5x-84 A is X^2 !!+!! 5x-84

OpenStudy (anonymous):

oh ok. I'm really bad a my positives an negatives.

OpenStudy (anonymous):

lol me too I have to recheck like 10 times for my answer

OpenStudy (anonymous):

I don't think its B ritz...

OpenStudy (anonymous):

Its not if it is I guess a bandit went and stole the 5x

OpenStudy (anonymous):

(x - 7)( x +12) = x^2 + 12x - 7x -84 = x^2 + 5x -84....so the answer is B